OpenStudy (anonymous):

Solve the polynomial equation by factoring and then using the zero-product principle: 4x^2(2x-3) -1(2x-3) = 0

OpenStudy (radar):

(2x-3)(2x-1)(2x+1) = 0 Double root x=3/2 x=1/2 x=-1/2

OpenStudy (anonymous):

oh i see I didnt factor 4x^2-1 completely thank you

OpenStudy (radar):

Good, I did not list every step, but you apparently followed it. Good luck with these.

OpenStudy (anonymous):

ya sometimes I just need to see it written out i guess thanks